OpenStudy (anonymous):

@jamesJ or whoever knws, limit x->0 (2sinx-1)?? dont knw where to begin on this one..

OpenStudy (asnaseer):

use the same method as James showed you in your last question.

OpenStudy (anonymous):

but what is sin? sin=1/cscx??

myininaya (myininaya):

f(x)=2 sin(x)-1 is continuous at x=0 therefore you just evaluate f(0) to find the limit here

OpenStudy (anonymous):

how can you tell it is continuous?

myininaya (myininaya):

because y=sin(x) is a continuous function It exists for any real input

OpenStudy (anonymous):

ok so on that note, is y=cosx also a continuous function?

myininaya (myininaya):

yep

myininaya (myininaya):

but y=tan(x) and y=sec(x) and y=cot(x) and y=csc(x) have discontinuities because at some point there bottom is zero

myininaya (myininaya):

like i'm saying there functions i just mentioned can be written in terms of cos(x) and sin(x) and remember a function is discontinuous( or one reason it can be discontinuous ) is if the bottom is zero

myininaya (myininaya):

these functions* as in y=tan(x),sec(x),cot(x),csc(x)

OpenStudy (anonymous):

u r a mind reader haha, that was exactly the question i was going to ask. thank u for ur help!!!!!